OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 1) and (1, 6)? A. 3/5 B. -5 C. -7 D. 5/3

OpenStudy (anonymous):

Please help I have no idea what the answer is

OpenStudy (anonymous):

Somebody please help

OpenStudy (ankit042):

slope = (y2-y1)/(x2-x1)

OpenStudy (tayy_lovee):

Have You Tired To Graph It Then Use Rise Over Run?

OpenStudy (anonymous):

yes but I got really confused and I didn't get any of the answers that are up there

OpenStudy (amistre64):

the slope formula is handy, but i tend to get the values to place in mixed around. i step it out as subtract one point from the other, and stack y/x for the slope

OpenStudy (amistre64):

(–2, 1 ) - ( 1 , 6) --------- x , y ; slope = y/x in simplest form

OpenStudy (anonymous):

how do you put it in simplest form

OpenStudy (anonymous):

is the answer a

OpenStudy (amistre64):

id say x=3, so 3 is on the bottom

OpenStudy (amistre64):

(–2, 1 ) - ( 1 , 6) --------- -3, -5 ; slope = 5/3 in simplest form (since - - = +)

OpenStudy (anonymous):

ohhh. so it is D ???

OpenStudy (tayy_lovee):